# Evaluating Language Model Math Reasoning via Grounding in Educational Curricula

### v1

@@ Abstract
Our work presents a novel angle for evaluating language models’ (LMs) mathematical abilities, by investigating whether they can discern skills and concepts enabled by math content. We contribute two datasets: one consisting of 385 fine-grained descriptions of K-12 math skills and concepts, or standards, from Achieve the Core (ATC [Uncaptioned image]), and another of 9.9K problems labeled with these standards (MathFish [Uncaptioned image]). Working with experienced teachers, we find that LMs struggle to tag and verify standards linked to problems, and instead predict labels that are close to ground truth, but differ in subtle ways. We also show that LMs often generate problems that do not fully align with standards described in prompts. Finally, we categorize problems in GSM8k using math standards, allowing us to better understand why some problems are more difficult to solve for models than others. [^1] Code, a data sample, and annotations can be found at https://anonymous.4open.science/r/math_anon-0720.
Evaluating Language Model Math Reasoning via
@@
Li Lucy1,2   Tal August1   Rose E. Wang3 Luca Soldaini1Courtney Allison4Kyle Lo1 1Allen Institute for AI   2University of California, Berkeley 3Stanford University  4EdReports lucy3_li@berkeley.edu
@@ 1 Introduction
As LMs are deployed in more real-world use cases, it is increasingly important to evaluate models with potential users. Throughout this paper, we work with curriculum specialists and teachers to center their voices when evaluating LM capabilities. Altogether, we hope that these datasets, along with the tooling we contribute for transforming data into task formats for LMs, can facilitate a more granular understanding of LM reasoning around skills and concepts in math content.
@@ 2 Related Work
Past work assessing mathematical reasoning capabilities of LMs have mainly focused on grade school arithmetic and algebra and problems that are easy to validate. These datasets may be written by annotators without formal pedagogical training (e.g. Cobbe et al., 2021; Patel et al., 2021; Amini et al., 2019; Ling et al., 2017; Miao et al., 2021), synthetically generated by LMs (Mitra et al., 2024; Liu et al., 2023a), and derived from advanced math competitions (Hendrycks et al., 2021). In contrast, our work characterizes real-world math curricula within an expansive, fine-grained taxonomy of math skills and concepts (§3). Concurrent work by Mishra et al. (2024) presents an approach for generating problems that align with math standards, but targets a limited set of easily validatable standards. Our work instead focuses on the task of labeling problems across 300+ comprehensive and challenging K-12 standards, and emphasizes ecological validity by working with educators (§4.2). Earlier work in education has identified math standards in problems with models such as SentenceBERT Li et al. (2024b), task-adapted BERT Shen et al. (2021), and support vector machines Karlovčec et al. (2012). Our work re-envisions this task into a format suitable for evaluating instruction-tuned LMs.
@@
LMs have increasing potential for improving math education through applications like automated feedback to educators (Jacobs et al., 2022; Demszky et al., 2023; Wang and Demszky, 2023), automated tutoring (Hobert and Meyer von Wolff, 2019; Wang et al., 2024b), and lesson planning (Malik et al., 2024; Kasneci et al., 2023). Evaluating models with education domain experts before real-world deployment is crucial to best align with educational needs and mitigate potential harms (Patikorn et al., 2019; Li et al., 2024b; Wang et al., 2024a). In our setting, we work with domain experts in a manner that centers real-world educational needs (§4.2).
@@ 3.1 Task Definition & Formats
In the initial phase of this work, we met with professional curriculum reviewers to better understand their process. We learned that identifying math standards in curricular materials is time-intensive, and reviewing a set of published materials can take 6-8 months to complete. This motivated us to investigate whether models can support reviewers by combing through subtle differences across standards. In collaboration with these curriculum reviewers, we identified two usage patterns in which these experts would employ language model assistance for performing curricular alignment—verifying whether a problem aligns to a given standard and tagging a problem with all aligned standards.
In the first, reviewers may wish to verify alignment, or check a problem against a single standard. For example, one may confirm publishers’ claims of alignment (§4.1), or evaluate whether generated materials follow $\mathcal{S}$ indicated in prompt instructions (§4.2). This task format is structured as a binary yes/no question (Figure 1): does a problem fully align with a given standard? Taking inspiration from textual entailment (Dagan et al., 2005) and claim verification (Thorne et al., 2018), this format also allows us to investigate models’ sensitivity to narrowing differences in positive and negative standards when we perturb them in queries.
Though verification is useful for checking problem alignment with a single standard, reviewers are also often tasked with identifying all aligned standards for a given problem. This task format, which we define as tagging, involves selecting from a provided set of candidates. Prompting language models with hundreds of candidate standards risks hitting context limits or models getting “lost in the middle” (Liu et al., 2023b). Instead, we take advantage of a domains-clusters-standards tree structure that experts use to organize standards into a hierarchy; related standards are grouped as leaves on this tree (Figure 1). We define the tagging task as follows: given a math problem, (1) start at the top of the tree and select the best domain/s $\mathcal{D}$ it teaches, (2) then traverse to each selected $\mathcal{D}$s’ subtrees, (3) repeat selection for cluster/s $\mathcal{C}$, (4) stop traversal upon reaching standards $\mathcal{S}$ at the leaves, and (5) tag the problem with selected $\mathcal{S}$. By traversing this tree, our tagging task format allows us to see whether models can make increasingly granular distinctions among adjacent concepts/skills.
@@
3.2 Math Standards & Organization
To facilitate these experiments, we also introduce two datasets: Achieve the Core (ATC [Uncaptioned image]) describes math standards and their organization, and MathFish [Uncaptioned image] links standard labels to problems (Table 1). Both involve Common Core State Standards (CCSS), which offer fine-grained and comprehensive coverage of K-12 math skills/concepts Porter et al. (2011).
ATC [Uncaptioned image] consists of CCSS standard labels (e.g. 4.NBT.A.1) and descriptions (e.g. Recognize that in a multi-digit whole number…) from Achieve the Core’s coherence map, which captures how pedagogical experts characterize, organize, and categorize math.333github.com/achievethecore/atc-coherence-map ATC [Uncaptioned image] includes two types of relationships among standards: as leaves in a topical hierarchy, or as a graph of conceptual connections. Our experiments on verifying (§4) and tagging (§5) use these relationships to show how related standards that differ subtly can challenge models.
3.3 K-12 Math Problems
We evaluate models’ abilities to identify math skills and concepts using publisher-labeled data pulled from curricular websites. Publicly accessible open educational resources (OER) provide a rich test bed for evaluating models, as they are designed to cover nearly all $\mathcal{S}$ within the grade levels they offer. We scrape pre-labeled problems from two OER that have been verified to be reputable by third party curriculum reviewers: Illustrative Mathematics and Fishtank Learning.555illustrativemathematics.org/math-curriculum and fishtanklearning.org/about Each problem (e.g. the Task in Figure 1) in this combined dataset, which we refer to as MathFish [Uncaptioned image], is a segment of these materials demarcated by $\mathcal{S}$ labels, which we map onto ATC [Uncaptioned image]’s natural language descriptions. A problem in MathFish [Uncaptioned image] can be labeled with multiple $\mathcal{S}$. Preprocessing details are in Appendix A.
@@ 3.4 Problems’ Alignment with Standards
Alignment is not always directly evident in problems’ language, and recognizing it requires a deeper understanding of cognitive processes. In Figure 1, each $\mathcal{C}$ and $\mathcal{S}$ relate to understanding the base ten number system, but differ in subtle ways. The purpose of the example problem on the left, based on the publisher’s description,666tasks.illustrativemathematics.org/content-standards/tasks/1562 is to teach students how an understanding of the place value system can facilitate division and multiplication by ten. If the problem were instead intended to enable students to Perform operations with multi-digit whole numbers and with decimals to hundredths, a wider range of whole numbers than powers of ten would have been included in the problem for students to fully practice those procedural skills. As another example of misalignment, the problem in Figure 1 does not align with the left $\mathcal{S}$ under Verification, as the problem involves place value understanding with decimals, which are not whole numbers. Thus, small differences in mathematical language have distinct consequences for student learning and assessment.
@@ 4.1 Can LMs discern (mis)aligned standards?
Error analysis.
We also qualitatively observe misconceptions around the solution strategy targeted by a problem. For example, in one problem, students are asked to Find the value of each expression mentally. 90-45, 270-45, 270-135, 360-135 … How did this observation—that the first numbers are all multiples of 90—help you find the value of the differences?777im.kendallhunt.com/k5/teachers/grade-4/unit-7/lesson-2/lesson In other words, this problem is designed to leverage multiplication to find numerical differences. GPT-4’s response when verifying a positive $\mathcal{S}$ related to multiplication instead claims that this problem focuses on subtraction and mental math strategies… rather than on multiplication of whole numbers... This example illustrates how identifying the concepts underlying problems requires recognizing how they are solved.
@@ 4.2 Study 1: Verifying standard alignment of LM-generated problems
Diliberti et al. (2024)’s survey of teachers showed that some of the most common uses of AI in classrooms include generating materials such as assessments, lesson plans, and assignments. In addition, problem generation is a common task in AI & education research (e.g. Norberg et al., 2023; Zhou et al., 2023; Wang et al., 2021; Mishra et al., 2024), though generations are rarely evaluated by in-domain experts. In this section, we investigate the utility of GPT-4, our best verifier from §4.1, by applying it onto assessing generated math problems, working alongside K-12 math teachers to also obtain their verification judgements.
@@ 8 Ethical Considerations
ATC [Uncaptioned image] is covered by a Creative Commons Public Domain Dedication License. Within MathFish [Uncaptioned image], Illustrative Mathematics is licensed as CC BY 4.0, while Fishtank Learning component is licensed under Creative Commons BY-NC-SA 4.0. Both sources are intended to be OER, which is defined as teaching, learning, and research materials that provides users free and perpetual permission to “retain, reuse, revise, remix, and redistribute” for educational purposes.888guides.library.columbia.edu/OER The transformed versions of these materials as datasets is licensed under ODC-By 1.0, and our code to reproduce our experiments is licensed under Apache 2.0.

### v3

# MathFish [Uncaptioned image]: Evaluating Language Model Math Reasoning via Grounding in Educational Curricula

@@ Abstract
To ensure that math curriculum is grade-appropriate and aligns with critical skills or concepts in accordance with educational standards, pedagogical experts can spend months carefully reviewing published math problems. Drawing inspiration from this process, our work presents a novel angle for evaluating language models’ (LMs) mathematical abilities, by investigating whether they can discern skills and concepts enabled by math content. We contribute two datasets: one consisting of 385 fine-grained descriptions of K-12 math skills and concepts, or standards, from Achieve the Core (ATC [Uncaptioned image]), and another of 9.9K math problems labeled with these standards (MathFish [Uncaptioned image]). We develop two tasks for evaluating LMs’ abilities to assess math problems: (1) verifying whether a problem aligns with a given standard, and (2) tagging a problem with all aligned standards. Working with experienced teachers, we find that LMs struggle to tag and verify standards linked to problems, and instead predict labels that are close to ground truth, but differ in subtle ways. We also show that LMs often generate problems that do not fully align with standards described in prompts, suggesting the need for careful scrutiny on use cases involving LMs for generating curricular materials. Finally, we categorize problems in GSM8k using math standards, allowing us to better understand why some problems are more difficult to solve for models than others.
MathFish [Uncaptioned image]: Evaluating Language Model Math Reasoning via
@@
Li Lucy1,2  Tal August1  Rose E. Wang3 Luca Soldaini1Courtney Allison4Kyle Lo1 1Allen Institute for AI  2University of California, Berkeley 3Stanford University  4EdReports lucy3_li@berkeley.edukylel@allenai.org
@@ 1 Introduction
As LMs are deployed in more real-world use cases, it is increasingly important to evaluate models with potential users. Throughout this paper, we work with curriculum specialists and teachers to center their voices when evaluating LM capabilities. These educational experts informed us that identifying math standards in curricular materials is time-intensive, and reviewing a set of published materials can take 6-8 months to complete. This motivates us to investigate whether models can support reviewers by combing through subtle differences across standards. Altogether, we hope that these datasets, along with the tooling we contribute for transforming data into task formats for LMs, can facilitate a more granular understanding of LM reasoning around skills and concepts in math content.
@@ 2 Related Work
Past work assessing mathematical reasoning capabilities of LMs have mainly focused on grade school arithmetic and algebra and problems that are easy to validate. These datasets may be written by annotators without formal pedagogical training (e.g. Cobbe et al., 2021; Patel et al., 2021; Amini et al., 2019; Ling et al., 2017; Miao et al., 2020), synthetically generated by LMs (Mitra et al., 2024; Liu et al., 2023a), and derived from advanced math competitions (Hendrycks et al., 2021). In contrast, our work characterizes real-world math curricula within an expansive, fine-grained taxonomy of math skills and concepts (§3). Concurrent work by Mishra et al. (2024) presents an approach for generating problems that align with math standards, but targets a limited set of easily validatable standards. Our work instead focuses on the task of labeling problems across 300+ comprehensive and challenging K-12 standards, and emphasizes ecological validity by working with educators (§4.2). Earlier work in education has identified math standards in problems with models such as SentenceBERT Li et al. (2024b), task-adapted BERT Shen et al. (2021), and support vector machines Karlovčec et al. (2012). Our work re-envisions this task into a format suitable for evaluating instruction-tuned LMs.
@@
LMs have increasing potential for improving math education through applications like automated feedback to educators (Jacobs et al., 2022; Demszky et al., 2023; Wang and Demszky, 2023), automated tutoring (Hobert and Meyer von Wolff, 2019; Wang et al., 2024), and lesson planning (Malik et al., 2024; Kasneci et al., 2023). Evaluating models with education domain experts before real-world deployment is crucial to best align with educational needs and mitigate potential harms (Patikorn et al., 2019; Li et al., 2024b; Lee et al., 2024). In our setting, we work with domain experts in a manner that centers real-world educational needs (§4.2).
@@ 3.1 Task Definition & Formats
What are task formats that both (1) evaluate LM ability to discern math skills or concepts in problems, as well as (2) reflect real-world usage patterns of LMs by pedagogical experts? In the initial phase of this work, we met with professional curriculum reviewers to better understand their processes and definitions for what does (or does not) constitute alignment to an educational standard. We identified two usage patterns in which these experts would employ language model assistance for performing curricular alignment—verifying whether a problem aligns to a given standard and tagging a problem with all aligned standards.
When verifying, we check a single problem against a single standard. This may arise when one is confirming publishers’ claims of alignment (§4.1), or when evaluating whether LM-generated materials follow the standard indicated in prompt instructions (§4.2). This task format is structured as a binary yes/no question (Figure 1): does a problem fully align with a given standard? Taking inspiration from textual entailment (Dagan et al., 2005) and claim verification (Thorne et al., 2018), this format also allows us to investigate models’ sensitivity to narrowing differences in positive and negative standards when we perturb them in queries.
Though verification is useful for checking problem alignment with a single standard, reviewers are also often tasked with identifying all aligned standards for a given problem. In tagging, we take a single problem and select all aligned standards from a provided set of candidates. Prompting language models with hundreds of candidate standards risks hitting context limits or models getting “lost in the middle” (Liu et al., 2023b). Instead, we take advantage of a domains-clusters-standards tree structure that experts use to organize standards into a hierarchy; related standards are grouped as leaves on this tree (Figure 1). We define the tagging task as follows: given a math problem, (1) start at the top of the tree and select the best domain/s $\mathcal{D}$ it teaches, (2) then traverse to each selected $\mathcal{D}$s’ subtrees, (3) repeat selection for cluster/s $\mathcal{C}$, (4) stop traversal upon reaching standards $\mathcal{S}$ at the leaves, and (5) tag the problem with selected $\mathcal{S}$. By traversing this tree, our tagging task format allows us to see whether models can make increasingly granular distinctions among adjacent concepts/skills.
@@
3.2 Math Standards & Organization
To facilitate these experiments, we introduce two datasets: Achieve the Core (ATC [Uncaptioned image]) describes math standards and their organization, and MathFish [Uncaptioned image] links standard labels to problems (Table 1). Both involve Common Core State Standards (CCSS), which offer fine-grained and comprehensive coverage of K-12 math skills/concepts Porter et al. (2011).
ATC [Uncaptioned image] consists of CCSS standard labels (e.g. 4.NBT.A.1) and descriptions (e.g. Recognize that in a multi-digit whole number…) from Achieve the Core’s coherence map, which captures how pedagogical experts characterize, organize, and categorize math.222github.com/achievethecore/atc-coherence-map ATC [Uncaptioned image] includes two types of relationships among standards: as leaves in a topical hierarchy, or as a graph of conceptual connections. Our experiments on verifying (§4) and tagging (§5) use these relationships to show how related standards that differ subtly can challenge models.
3.3 K-12 Math Problems
We evaluate models’ abilities to identify math skills and concepts using publisher-labeled data pulled from curricular websites. Publicly accessible open educational resources (OER) provide a rich test bed for evaluating models, as they are designed to cover nearly all $\mathcal{S}$ within the grade levels they offer. We scrape pre-labeled problems from two OER that have been verified to be reputable by third party curriculum reviewers: Illustrative Mathematics and Fishtank Learning.444illustrativemathematics.org/math-curriculum and fishtanklearning.org/about Each problem (e.g. the Task in Figure 1) in this combined dataset, which we refer to as MathFish [Uncaptioned image], is a segment of these materials demarcated by $\mathcal{S}$ labels, which we map onto ATC [Uncaptioned image]’s natural language descriptions. A problem in MathFish [Uncaptioned image] can be labeled with multiple $\mathcal{S}$. Preprocessing details are in Appendix A.
@@ 3.4 Problems’ Alignment with Standards
Alignment is not always directly evident in problems’ language, and recognizing it requires a deeper understanding of cognitive processes. In Figure 1, each $\mathcal{C}$ and $\mathcal{S}$ relate to understanding the base ten number system, but differ in subtle ways. The purpose of the example problem on the left, based on the publisher’s description,555tasks.illustrativemathematics.org/content-standards/tasks/1562 is to teach students how an understanding of the place value system can facilitate division and multiplication by ten. If the problem were instead intended to enable students to Perform operations with multi-digit whole numbers and with decimals to hundredths, a wider range of whole numbers than powers of ten would have been included in the problem for students to fully practice those procedural skills. As another example of misalignment, the problem in Figure 1 does not align with the left $\mathcal{S}$ under Verification, as the problem involves place value understanding with decimals, which are not whole numbers. Thus, small differences in mathematical language have distinct consequences for student learning and assessment.
@@ 4.1 Can LMs discern (mis)aligned standards?
Error analysis.
We also qualitatively observe misconceptions around the solution strategy targeted by a problem. For example, in one problem, students are asked to Find the value of each expression mentally. 90-45, 270-45, 270-135, 360-135 … How did this observation—that the first numbers are all multiples of 90—help you find the value of the differences?666im.kendallhunt.com/k5/teachers/grade-4/unit-7/lesson-2/lesson In other words, this problem is designed to leverage multiplication to find numerical differences. GPT-4’s response when verifying a positive $\mathcal{S}$ related to multiplication instead claims that this problem focuses on subtraction and mental math strategies… rather than on multiplication of whole numbers... This example illustrates how identifying the concepts underlying problems requires recognizing how they are solved.
@@ 4.2 Study 1: Verifying standard alignment of LM-generated problems
Diliberti et al. (2024)’s survey of teachers showed that some of the most common uses of AI in classrooms include generating materials such as assessments, lesson plans, and assignments. In addition, problem generation is a common task in AI & education research (e.g. Norberg et al., 2023; Zhou et al., 2023; Wang et al., 2021; Mishra et al., 2024; Shah et al., 2024), though generations are rarely evaluated by in-domain experts. In this section, we investigate the utility of GPT-4, our best verifier from §4.1, by applying it onto assessing generated math problems, working alongside K-12 math teachers to also obtain their verification judgements.
@@ 8 Ethical Considerations
ATC [Uncaptioned image] is covered by a Creative Commons Public Domain Dedication License. Within MathFish [Uncaptioned image], Illustrative Mathematics is licensed as CC BY 4.0, while Fishtank Learning component is licensed under Creative Commons BY-NC-SA 4.0. Both sources are intended to be OER, which is defined as teaching, learning, and research materials that provides users free and perpetual permission to “retain, reuse, revise, remix, and redistribute” for educational purposes.777guides.library.columbia.edu/OER The transformed versions of these materials as datasets is licensed under ODC-By 1.0, and our code to reproduce our experiments is licensed under Apache 2.0.
9 Acknowledgements
We thank the sixteen teachers from EdReports not only for their annotations, but also for their conversations which greatly informed our work. We also thank curriculum specialists at EdReports for patiently guiding us through the ecosystem of “real-world” math curriculum and curriculum review. Our work was generously funded by the Bill & Melinda Gates Foundation.